OpenStudy (anonymous):

Factorise the following expressions; a) x²-x)-(2-2x) b) 4x²(x-2)+2x(x-2)

OpenStudy (anonymous):

Go ahead and start factorizing it. then tell us where you're stuck

OpenStudy (lgbasallote):

hint: \[(2 - 2x) = -(2x - 2)\]

OpenStudy (callisto):

(x²-x)-(2-2x) = x² -x -2 + 2x = x² +x -2 = x² +2x -x -2 = x(x+2) - (x+2) = (x-1)(x+2)

OpenStudy (lgbasallote):

OR \[(x^2 - x) + (2x - 2)\] \[x(x - 1) + 2(x-1)\] \[(x-1)(x+2)\] whichever you like better

OpenStudy (callisto):

4x²(x-2)+2x(x-2) = (4x²+2x)(x-2) Since there is 2x in the terms in the first bracket, take it out = 2x(2x+1)(x-2)

OpenStudy (callisto):

My favourite way to do (a): (x²-x)-(2-2x) = x² -x -2 + 2x = x² +x -2  = (x+2)(x-1)